# UVE: Are MLLMs Unified Evaluators for AI-Generated Videos?

### v2

@@ 4.4 Data Annotation and Quality Review
With the collected videos, the authors of this paper, who have rich research experience in VGMs and MLLMs and are proficient in English, annotate the evaluation aspects and pairwise video preferences. Specifically, videos generated by the same text prompt are presented in a pairwise manner to the annotators, who are asked to determine a preference choice from $\mathbf{O}$ and identify the aspect that influenced their decision. To minimize subjectivity, annotators are instructed to only annotate video pairs for which they could confidently make a preference judgment, and the remaining video pairs are discarded. Following this procedure, we obtained 4,042 pairwise preference annotations and retained 1,230 individual videos. The distributions of the videos and preference annotations are illustrated in Figure 1.
@@
Dataset Aspects Eval Task Latest VGM FETV (Liu et al., 2023b) 9 Single ZeroScope (2023.06) VBench (Huang et al., 2024) 16 Pair VideoCrafter1 (2023.10) EvalCrafter (Liu et al., 2024b) 4 Single Gen-2 (2023.12) T2VQA-DB (Kou et al., 2024) 1 Single AnimateDiff (2023.12) VideoFeedback (He et al., 2024) 5 Single SORA (2024.02) LGVQ (Zhang et al., 2024d) 3 Single Gen-2 (2023.12) TVGE (Wu et al., 2024c) 2 Single Gen-2 (2023.12) GAIA (Mialon et al., 2023) 3 Single Mora (2024.03) Q-Bench-Video (Zhang et al., 2024c) 4 Understanding SORA (2024.02) AIGV-Assessor (Wang et al., 2024a) 4 Single&Pair SORA (2024.02) UVE-Bench (Ours) 15 Single&Pair HunyuanVideo (2024.12)
@@ 4.5 Comparison with Existing AIGV Datasets
Table 3 compares UVE-Bench with existing AIGV datasets across three dimensions. First, current datasets typically focus on a few basic aspects (such as video-text alignment, static and temporal visual quality), making them not suitable for assessing MLLMs as unified evaluators across various aspects. Second, existing datasets predominantly provide human annotations in the form of single-video rating, which not only precludes video pair comparison analysis but also introduces greater subjectivity compared to pairwise preference annotation (Wu et al., 2024b; Zhang et al., 2024e). Third, the VGMs employed in existing datasets are relatively outdated, failing to represent the current SOTA performance of AIGVs. Identifying the weaknesses of such videos is both important and challenging for automatic AIGV evaluators. UVE-Bench addresses these limitations and provides a more comprehensive and accurate assessment for automatic AIGV evaluation models.
Method Model Size Overall Dynamic Subject Motion Camera Motion Light& Color Overall Static Technical Quality Structural Correctness Aesthetic Quality Overall Temporal Appearance Consistency Motion Naturalness Temporal Flickering Overall Alignment Motion Alignment Appearance Alignment AVG Random - 48.8 49.5 48.9 49.8 49.2 47.6 49.0 48.1 49.0 48.6 48.2 46.2 48.4 48.6 48.3 48.6 Video-LLaVA 7B 52.4 74.8 63.4 47.6 47.8 66.1 44.5 63.4 44.1 51.3 55.5 48.6 59.4 54.9 66.8 54.5 LongVA-DPO 7B 59.8 76.2 69.9 57.5 62.4 74.9 56.2 72.0 56.0 47.3 40.7 54.3 68.7 63.9 71.6 61.6 ShareGPT4Video 8B 77.5 81.0 77.1 82.5 59.4 68.7 54.1 69.4 54.4 48.1 42.6 60.9 62.7 54.3 70.3 63.9 VideoLLaMA2.1 7B 72.1 80.5 67.1 77.2 61.4 78.7 47.5 72.6 46.1 50.9 50.2 61.8 72.6 65.7 77.7 64.4 mPLUG-Owl3 7B 78.6 84.8 77.8 79.9 76.0 83.1 55.6 80.0 59.8 59.5 42.0 72.0 80.4 75.2 87.6 72.6 VideoChat2-Mistral 7B 83.1 92.2 89.6 74.9 68.1 76.2 53.8 74.1 58.7 58.3 52.8 85.6 75.6 78.0 80.9 72.6 MiniCPM-V-2.6 8B 81.4 86.3 80.3 88.8 70.9 75.0 52.9 80.6 61.1 59.4 51.7 70.9 82.1 74.3 90.8 73.4 LLaVA-OneVision 7B 81.0 87.6 83.2 84.9 70.7 78.2 50.6 83.1 62.9 60.4 41.5 85.9 79.3 66.9 86.7 73.0 LLaVA-OneVision 72B 82.3 87.8 78.4 88.2 71.3 77.6 60.2 81.5 64.6 61.9 39.9 86.8 84.4 71.7 93.5 75.0 LLaVA-Video 7B 80.4 85.5 80.9 81.5 66.2 74.2 49.5 75.7 58.3 58.2 39.3 82.3 80.5 69.9 90.0 71.0 LLaVA-Video 72B 82.8 86.1 82.9 86.9 70.2 80.0 55.4 77.7 60.1 59.2 40.4 83.5 84.8 73.7 94.6 74.0 Qwen2-VL 7B 84.6 89.7 94.2 79.7 64.6 67.3 50.7 70.6 51.1 51.3 48.0 62.7 85.4 78.7 92.2 70.9 Qwen2-VL 72B 86.5 92.6 92.7 86.0 70.6 76.9 60.2 83.4 52.5 58.0 48.9 71.0 89.0 81.8 95.0 75.4 InternVL-2.5-MPO 8B 81.3 86.1 80.4 88.0 68.1 77.9 53.6 77.5 60.9 54.9 50.9 72.5 80.6 73.2 90.5 72.6 InternVL-2.5-MPO 78B 84.3 86.6 82.4 91.6 72.8 84.0 67.2 82.3 61.5 65.2 61.8 88.4 87.4 79.3 95.5 78.2 GPT-4o - 79.0 84.3 74.9 81.8 74.0 81.6 65.2 84.2 70.0 79.8 54.0 58.6 80.8 77.2 89.6 75.7 VideoScore-v1.1 8B 57.4 - - - 40.1 30.4 47.7 41.9 - 43.1 36.1 - 38.9 - - - VBench - 87.8 - - - - 62.5 - 75.6 - 54.0 50.2 - 54.4 - - -
Table 4: Performance of single video rating measured by $\mathcal{A}^{\text{single}}$ in Eq. 3. The best and second-best results are highlighted with yellow and light yellow, respectively.
Method Model Size Overall Dynamic Subject Motion Camera Motion Light& Color Overall Static Technical Quality Structural Correctness Aesthetic Quality Overall Temporal Appearance Consistency Motion Naturalness Temporal Flickering Overall Alignment Motion Alignment Appearance Alignment AVG Random - 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 LLaVA-OneVision 7B 44.2 52.6 40.0 38.5 35.8 35.2 25.7 51.0 37.4 23.9 29.1 52.0 43.1 39.7 43.4 38.6 LLaVA-OneVision 72B 36.5 55.1 40.0 53.8 37.0 51.4 22.2 54.1 25.6 43.3 34.5 45.7 63.0 58.9 69.7 44.3 LLaVA-Video 7B 37.0 52.6 37.5 44.9 44.4 43.8 38.9 62.2 33.7 28.9 31.1 57.5 43.1 39.7 46.9 41.1 LLaVA-Video 72B 42.5 57.7 32.5 56.4 41.6 55.2 31.2 60.2 32.6 41.1 27.7 55.9 60.3 53.0 66.2 46.1 Qwen2-VL 7B 46.4 56.4 43.8 39.7 42.8 38.1 20.8 54.1 29.8 24.4 29.7 42.5 54.9 50.3 57.9 41.1 Qwen2-VL 72B 51.4 66.7 71.2 53.8 47.7 62.9 19.4 60.2 41.0 40.0 31.8 40.2 69.7 62.9 78.6 51.6 InternVL-2.5-MPO 8B 42.5 51.3 33.8 43.6 38.3 31.4 40.3 55.1 35.1 32.2 27.0 44.9 53.5 50.3 53.8 41.8 InternVL-2.5-MPO 78B 43.1 57.7 41.2 61.5 54.7 62.9 27.1 71.4 45.2 47.8 33.8 70.9 67.7 55.6 76.6 53.7 GPT-4o - 42.0 48.7 38.8 59.0 53.5 54.3 41.0 71.4 44.9 38.9 31.8 59.8 58.9 57.0 61.4 50.2 Human - 87.3 85.3 87.3 85.3 90.0 92.0 88.0 91.3 88.0 90.0 78.7 92.0 88.0 84.7 92.0 88.0
(a) Single Video Rating.
(b) Video Pair Comparison.
Figure 2: Results of different prompting strategies.
Figure 3: Results of different scoring strategies with Qwen2-VL-7B for single video rating.
@@ 5.1 Experimental Settings
We conduct zero-shot evaluations on 16 MLLMs, including 15 open-sourced models along with the proprietary GPT-4o model (OpenAI, 2024). Additionally, we assess two evaluation systems specialized in particular aspects: VBench (Huang et al., 2024) employs off-the-shelf models for evaluation and VideoScore (He et al., 2024) is a fine-tuned version of Mantis-Idefics2-8B (Jiang et al., 2024), trained to imitate human rating scores. Details of the automatic evaluators are presented in Appendix A.1. To establish a human baseline, we engage three annotators to perform video pair comparisons on a subset of UVE-Bench, evaluating 50 samples for each subaspect.
@@
In this evaluation mode, we utilize 12 frames per video. As shown in Figure 1 (c), the original UVE-Bench contains more “$\mathcal{V}_{1/2}\text{better}$” compared to “same good/bad”. While this imbalance does not compromise the fairness of single video ratings (since $\mathcal{A}^{\text{single}}$ handles “better” and “same” independently) it introduces a bias toward “$\mathcal{V}_{1/2}\text{better}$” when computing four-way selection accuracy for video pair comparisons. To mitigate this bias, we create a subset of 2,411 samples, with a more balanced distribution across the four preference categories, in evaluating video pair comparison.
(a) Single Video Rating.
(b) Video Pair Comparison.
Figure 4: Results with varying numbers of video frames. When the frame number equals to 0, MLLMs receive pure-text prompt without visual input.
Figure 5: Results of direct video pair comparison versus adapting single video rating to pairwise comparison.
@@ 5.2 Performance of MLLMs as Unified Evaluator
Table 4 presents the results of MLLMs in single video rating, from which we can obtain the following findings: (1) SOTA MLLMs demonstrate strong capabilities in assessing Dynamic Degree and Video-Text Alignment, with Qwen2-VL-72B achieving over $80$ $\mathcal{A}^{\text{single}}$ score across all eight subaspects. (2) current MLLMs show limitations in evaluating Temporal Quality aspects, which require a nuanced understanding of video temporal dynamics. Additionally, their performance in Motion Alignment lags behind that of Appearance Alignment. These observations echo with previous research (Liu et al., 2024c; Shangguan et al., 2024) which reveal the limitation of MLLMs in fine-grained video temporal understanding. Interestingly, Video-LLaVA and VideoChat2-Mistral, despite their lower average $\mathcal{A}^{\text{single}}$, perform well in assessing motion naturalness, surpassing the 72B Qwen2-VL and LLaVA-OneVision. We attribute this to their use of native video encoders, rather than the frame-by-frame processing approach used by most MLLMs. (3) While advanced MLLMs can effectively assess Technical Quality and Aesthetic Quality, they still struggle to identify incorrect subject structures (e.g., human hands with six fingers). (4) MLLMs’ performance in single video rating strongly correlates with their general multimodal understanding capability: Larger models, such as Qwen2-VL-72B, InternVL2.5-78B-MPO, and GPT-4o, demonstrate superior performance compared to the smaller ones at 7B scale. (5) The unified evaluators consistently outperform or match specialized methods across all evaluation aspects. This indicates that utilizing MLLMs for unified AIGV evaluation is a promising direction, which showcases both versatility and effectiveness.
@@
As shown in Table 5, the MLLMs’ performance in video pair comparison exhibits a similar pattern as single video rating: They achieve relative higher accuracy when evaluating Dynamic Degree subaspects, Technical Quality, Aesthetic Quality and Appearance Alignment, while showing weakness in assessing Structural Correctness, Motion Naturalness and Motion Alignment. Additionally, the significant performance gap of over 30 points between SOTA MLLMs and human evaluators indicates that current MLLMs have not yet achieved reliable pairwise video quality assessment capabilities.
5.3 Effect of Prompting Strategy
We investigate how different prompting strategies affect model performance by modifying our original prompt templates (Table 1). We test two variations: (1) removing the aspect-specific descriptions while maintaining only the question and answer instructions, and (2) for video pair comparisons, eliminating video order indicators (“The first/second video”). The latter modification was motivated by Zhang et al. (2024f), which emphasizes the importance of specifying relative order in image quality comparisons. The modified prompts can be found in the supplementary material.

### v4

@@ 3.4 Data Annotation and Quality Review
With the collected videos, the authors of this paper, who have rich research experience in VGMs and MLLMs and are proficient in English, annotate the evaluation aspects and pairwise video preferences. Specifically, videos generated by the same text prompt are presented in a pairwise manner to the annotators, who are asked to determine a preference choice from $\mathbf{O}$ and identify the aspect that influenced their decision. To minimize subjectivity, annotators are instructed to only annotate video pairs for which they could confidently make a preference judgment, and the remaining video pairs are discarded. Following this procedure, we obtained 4,042 pairwise preference annotations and retained 1,230 individual videos. The distributions of the videos and preference annotations are illustrated in Fig. 2.
@@
Dataset Aspects Eval Task Latest VGM FETV [39] 9 Single ZeroScope (2023.06) VBench [21] 16 Pair VideoCrafter1 (2023.10) EvalCrafter [40] 4 Single Gen-2 (2023.12) T2VQA-DB [26] 1 Single AnimateDiff (2023.12) VideoFeedback [17] 5 Single SORA (2024.02) LGVQ [75] 3 Single Gen-2 (2023.12) TVGE [63] 2 Single Gen-2 (2023.12) GAIA [43] 3 Single Mora (2024.03) Q-Bench-Video [74] 4 Understanding SORA (2024.02) AIGV-Assessor [56] 4 Single&Pair SORA (2024.02) UVE-Bench (Ours) 15 Single&Pair HunyuanVideo (2024.12)
@@ 3.5 Comparison with Existing AIGV Datasets
Tab. 3 compares UVE-Bench with existing AIGV datasets across three dimensions. First, current datasets typically focus on a few basic aspects (such as video-text alignment, static and temporal visual quality), making them not suitable for assessing MLLMs as unified evaluators across various aspects. Second, existing datasets predominantly provide human annotations in the form of single-video rating, which not only precludes video pair comparison analysis but also introduces greater subjectivity compared to pairwise preference annotation [62, 76]. Third, the VGMs employed in existing datasets are relatively outdated, failing to represent the current SOTA performance of AIGVs. Identifying the weaknesses of such videos is both important and challenging for automatic AIGV evaluators. UVE-Bench addresses these limitations and provides a more comprehensive and accurate assessment for automatic AIGV evaluation models.
@@ 4.1 Experimental Settings
We conduct zero-shot evaluations on 18 MLLMs, including 15 open-sourced models along with three proprietary models: GPT-4o [45], Seed1.5-VL [15] and Gemini2.5-Flash [14]. Additionally, we assess five evaluation systems specialized in particular aspects: VBench [21], VideoScore [17], UMTScore [39], VIDEOCON-PHYSICS [2] and DOVER [60]. Details of these automatic evaluators are presented in Appendix C.1. To establish a human baseline, we engage three annotators to perform video pair comparisons on a subset of UVE-Bench, evaluating 50 samples for each subaspect.
@@
In this evaluation mode, we utilize 12 frames per video. As shown in Fig. 2 (c), the original UVE-Bench contains more $\mathcal{V}_{1/2}\text{better}$” compared to “same good/bad”. While this imbalance does not compromise the fairness of single video ratings (since $\mathcal{A}^{\text{single}}$ handles “better” and “same” independently) it introduces a bias toward $\mathcal{V}_{1/2}\text{better}$” when computing four-way selection accuracy for video pair comparisons. To mitigate this bias, we create a subset of 2,411 samples, with a more balanced distribution across the four preference categories, in evaluating video pair comparison.
Table 4: Performance of single video rating measured by $\mathcal{A}^{\text{single}}$ in Eq. 3. The best and second-best results are highlighted with yellow and light yellow, respectively. Abbreviations: SM (subject motion), CM (camera motion), LC (light&color), TQ (technical quality), SC (structural correctness), AQ (aesthetic quality), AC (appearance consistency), MN (motion naturalness), TF (temporal flickering), MA (motion alignment), AA (appearance alignment).
Method Model Size Overall Dynamic SM CM LC Overall Static TQ SC AQ Overall Temporal AC MN TF Overall Alignment MA AA AVG Random - 48.8 49.5 48.9 49.8 49.2 47.6 49.0 48.1 49.0 48.6 48.2 46.2 48.4 48.6 48.3 48.6 Specialized Evaluators VideoScore-v1.1 8B 57.4 - - - 40.1 30.4 47.7 41.9 - 43.1 36.1 - 38.9 - - - VBench - 87.8 - - - - 62.5 - 75.6 - 54.0 50.2 - 54.4 - - - UMTScore - - - - - - - - - - - - - 66.4 - - - VIDEOCON-PHYSICS 7B - - - - - - - - - - 53.4 - 68.7 - - - DOVER 58M - - - - - 69.2 - 80.3 - - - - - - - - Unified Evaluators Video-LLaVA 7B 52.4 74.8 63.4 47.6 47.8 66.1 44.5 63.4 44.1 51.3 55.5 48.6 59.4 54.9 66.8 54.5 LongVA-DPO 7B 59.8 76.2 69.9 57.5 62.4 74.9 56.2 72.0 56.0 47.3 40.7 54.3 68.7 63.9 71.6 61.6 ShareGPT4Video 8B 77.5 81.0 77.1 82.5 59.4 68.7 54.1 69.4 54.4 48.1 42.6 60.9 62.7 54.3 70.3 63.9 VideoLLaMA2.1 7B 72.1 80.5 67.1 77.2 61.4 78.7 47.5 72.6 46.1 50.9 50.2 61.8 72.6 65.7 77.7 64.4 mPLUG-Owl3 7B 78.6 84.8 77.8 79.9 76.0 83.1 55.6 80.0 59.8 59.5 42.0 72.0 80.4 75.2 87.6 72.6 VideoChat2-Mistral 7B 83.1 92.2 89.6 74.9 68.1 76.2 53.8 74.1 58.7 58.3 52.8 85.6 75.6 78.0 80.9 72.6 MiniCPM-V-2.6 8B 81.4 86.3 80.3 88.8 70.9 75.0 52.9 80.6 61.1 59.4 51.7 70.9 82.1 74.3 90.8 73.4 LLaVA-OneVision 7B 81.0 87.6 83.2 84.9 70.7 78.2 50.6 83.1 62.9 60.4 41.5 85.9 79.3 66.9 86.7 73.0 LLaVA-OneVision 72B 82.3 87.8 78.4 88.2 71.3 77.6 60.2 81.5 64.6 61.9 39.9 86.8 84.4 71.7 93.5 75.0 LLaVA-Video 7B 80.4 85.5 80.9 81.5 66.2 74.2 49.5 75.7 58.3 58.2 39.3 82.3 80.5 69.9 90.0 71.0 LLaVA-Video 72B 82.8 86.1 82.9 86.9 70.2 80.0 55.4 77.7 60.1 59.2 40.4 83.5 84.8 73.7 94.6 74.0 Qwen2-VL 7B 84.6 89.7 94.2 79.7 64.6 67.3 50.7 70.6 51.1 51.3 48.0 62.7 85.4 78.7 92.2 70.9 Qwen2-VL 72B 86.5 92.6 92.7 86.0 70.6 76.9 60.2 83.4 52.5 58.0 48.9 71.0 89.0 81.8 95.0 75.4 InternVL-2.5-MPO 8B 81.3 86.1 80.4 88.0 68.1 77.9 53.6 77.5 60.9 54.9 50.9 72.5 80.6 73.2 90.5 72.6 InternVL-2.5-MPO 78B 84.3 86.6 82.4 91.6 72.8 84.0 67.2 82.3 61.5 65.2 61.8 88.4 87.4 79.3 95.5 78.2 GPT-4o - 79.0 84.3 74.9 81.8 74.0 81.6 65.2 84.2 70.0 79.8 54.0 58.6 80.8 77.2 89.6 75.7 Seed1.5-VL 20B Act. 83.2 91.2 83.7 89.5 82.4 82.9 66.8 88.8 70.5 78.6 59.5 70.1 84.2 79.0 94.2 80.0
Method Model Size Overall Dynamic SM CM LC Overall Static TQ SC AQ Overall Temporal AC MN TF Overall Alignment MA AA AVG Random - 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 Unified Evaluators LLaVA-OneVision 7B 44.2 52.6 40.0 38.5 35.8 35.2 25.7 51.0 37.4 23.9 29.1 52.0 43.1 39.7 43.4 38.6 LLaVA-OneVision 72B 36.5 55.1 40.0 53.8 37.0 51.4 22.2 54.1 25.6 43.3 34.5 45.7 63.0 58.9 69.7 44.3 LLaVA-Video 7B 37.0 52.6 37.5 44.9 44.4 43.8 38.9 62.2 33.7 28.9 31.1 57.5 43.1 39.7 46.9 41.1 LLaVA-Video 72B 42.5 57.7 32.5 56.4 41.6 55.2 31.2 60.2 32.6 41.1 27.7 55.9 60.3 53.0 66.2 46.1 Qwen2-VL 7B 46.4 56.4 43.8 39.7 42.8 38.1 20.8 54.1 29.8 24.4 29.7 42.5 54.9 50.3 57.9 41.1 Qwen2-VL 72B 51.4 66.7 71.2 53.8 47.7 62.9 19.4 60.2 41.0 40.0 31.8 40.2 69.7 62.9 78.6 51.6 InternVL-2.5-MPO 8B 42.5 51.3 33.8 43.6 38.3 31.4 40.3 55.1 35.1 32.2 27.0 44.9 53.5 50.3 53.8 41.8 InternVL-2.5-MPO 78B 43.1 57.7 41.2 61.5 54.7 62.9 27.1 71.4 45.2 47.8 33.8 70.9 67.7 55.6 76.6 53.7 GPT-4o - 42.0 48.7 38.8 59.0 53.5 54.3 41.0 71.4 44.9 38.9 31.8 59.8 58.9 57.0 61.4 50.2 Seed1.5-VL 20B Act. 52.5 56.4 47.5 61.5 51.0 52.4 44.4 62.2 48.6 36.7 35.8 65.4 56.9 53.6 61.4 51.6 Gemini2.5-Flash - 55.8 60.3 45.0 59.0 55.6 50.5 43.1 64.3 50.8 41.7 38.5 60.6 65.0 62.3 66.2 54.6 Human - 87.3 85.3 87.3 85.3 90.0 92.0 88.0 91.3 88.0 90.0 78.7 92.0 88.0 84.7 92.0 88.0
(a) Results of different prompting strategies.
(b) Results of different scoring strategies with Qwen2-VL-7B for single video rating.
@@ 4.2 Performance of MLLMs as Unified Evaluator
Tab. 4 presents the results of MLLMs in single video rating, from which we can obtain the following findings: (1) SOTA MLLMs demonstrate strong capabilities in assessing Dynamic Degree and Video-Text Alignment, with Qwen2-VL-72B achieving over 8080 $\mathcal{A}^{\text{single}}$ score across all eight subaspects. (2) current MLLMs show limitations in evaluating Temporal Quality aspects, which require a nuanced understanding of video temporal dynamics. Additionally, their performance in Motion Alignment lags behind that of Appearance Alignment. These observations echo with previous research [41, 52] which reveal the limitation of MLLMs in fine-grained video temporal understanding. Interestingly, Video-LLaVA and VideoChat2-Mistral, despite their lower average $\mathcal{A}^{\text{single}}$, perform well in assessing motion naturalness, surpassing the 72B Qwen2-VL and LLaVA-OneVision. We attribute this to their use of native video encoders, rather than the frame-by-frame processing approach used by most MLLMs. (3) While advanced MLLMs can effectively assess Technical Quality and Aesthetic Quality, they still struggle to identify incorrect subject structures (e.g., human hands with six fingers). (4) MLLMs’ performance in single video rating strongly correlates with their general multimodal understanding capability: Larger models, such as Qwen2-VL-72B, InternVL2.5-78B-MPO, GPT-4o and Seed1.5-VL demonstrate superior performance compared to the smaller ones at 7B scale. (5) The unified evaluators consistently outperform or match specialized methods across all evaluation aspects. This indicates that utilizing MLLMs for unified AIGV evaluation is a promising direction, which showcases both versatility and effectiveness.
@@
As shown in Tab. 5, the MLLMs’ performance in video pair comparison exhibits a similar pattern as single video rating: They achieve relative higher accuracy when evaluating Dynamic Degree subaspects, Technical Quality, Aesthetic Quality and Appearance Alignment, while showing weakness in assessing Structural Correctness, Motion Naturalness and Motion Alignment. Additionally, the significant performance gap of over 30 points between SOTA MLLMs and human evaluators indicates that current MLLMs have not yet achieved reliable pairwise video quality assessment capabilities.
4.3 Factors Influencing Unified Evaluator Performance
4.3.1 Impact of Prompting Strategy
We investigate how different prompting strategies affect model performance by modifying our original prompt templates (Tab. 2.1). We test two variations: (1) removing the aspect-specific descriptions while maintaining only the question and answer instructions, and (2) for video pair comparisons, eliminating video order indicators (“The first/second video”). The latter modification was motivated by Zhang et al. [77], which emphasizes the importance of specifying relative order in image quality comparisons. The modified prompts can be found in the supplementary material.